The sun has a radiated power of 3.9 · 1026 W, a surface temperature of 5800 K and a radius of 6.96 · 108 m. Examine whether the sun radiates like a black body?
1
Expert's answer
2021-05-17T16:24:57-0400

Radiation energy Q emitted by an object in time t, object has surface area A, emissivity e and kelvin temperature T is:

"Q=e \sigma T^4At \\

\sigma = 5.87 \times 10^{-8} \;J/(s \times m^2 \times T^4) \\

Q = 3.9 \times 10^{26} \;W \\

r = 6.96 \times 10^8 \;m \\

T = 5800 =5.8 \times 10^3\;K \\

t = 1 \;s"

The sun is an spherical shape object. The surface area of the sun is

"A = 4 \pi r^2 \\

= 4 \pi (6.96 \times 10^8)^2 \\

= 6.087 \times 10^{18} \;m^2 \\

e = \frac{Q}{\sigma T^4At} \\

= \frac{3.9 \times 10^{26}}{5.87 \times 10^{-8} \times (5.8 \times 10^3)^4 \times 6.087 \times 10^{18} \times 1} \\

= \frac{3.9 \times 10^{26}}{4.04 \times 10^{26}} \\

= 0.965 ≈ 1"

e = 1 for black body

So, the sun radiates like a black body.